OpenStudy (anonymous):

I'm stuck on this problem... Simplify: square root of 2/3 times square root 3/5 Show steps so I can follow along, please?

OpenStudy (anonymous):

\[\sqrt{\frac{2}{3}}\times \sqrt{\frac{3}{5}}=\sqrt{\frac{2\times 3}{3\times 5}}=\sqrt{\frac{2}{5}}\]

OpenStudy (anonymous):

Ah! I see where I was going wrong... I didn't simplify square root of 6/square root of 15. Thank you!!